# Predictive Coding Networks and Inference Learning: Tutorial and Survey

### v1

@@ A.2.1. General Idea
$\equiv\int\prod_{\ell=0}^{L}d\bm{z}^{\ell}q(\{\bm{z}^{\ell}\})\ln\frac{q(\{\bm{z}^{\ell}\})}{p_{\theta}(\bm{x},\{\bm{z}^{\ell}\})}$
$=\underbrace{\int\prod_{\ell=0}^{L}\mathrm{d}\bm{z}^{\ell}q(\bar{\bm{z}})\ln q(\bar{\bm{z}})}_{H[\bar{q}]}-\underbrace{\int\prod_{\ell=0}^{L}\mathrm{d}\bm{z}^{\ell}q(\bar{\bm{z}})E(\bm{x},\bar{\bm{z}})}_{E_{\bar{q}}}~{},$
@@ A.2.2. PCNs of Rao & Ballard: Delta-function variational
Recent work (millidge_predictive_2022-1, ) presents an attempted derivation of PC using a delta function as the variational posterior, in which it is claimed that the VFE simply becomes equal to the energy. Unfortunately however, this derivation contains a basic mathematical error, which we explain below. An alternative derivation (buckley_free_2017, ) also exists that instead uses a Gaussian as variational posterior, approximated using a Taylor series. Under certain assumptions, this leads to the claimed result. We review the both here and in the next subsection.
@@
In (millidge_predictive_2022, ) this is erroneously taken to be zero. We speculate that this is based on the intuition that the “entropy” of a deterministic function like $\delta(z)$ should vanish. In fact however, the differential entropy of the Dirac delta function is $\infty$- ∞. If however one treats this as an additive constant to be dropped, then we have only the $E_{q}$ term, so that
@@
For this reason, (millidge_predictive_2022, ) identify the VFE as simply the energy (or negative complete data log-likelihood) in this case.
@@ A.2.3. PCNs of Rao & Ballard: Gaussian variational
Now we instead take $q$ to be a multivariate Gaussian centered around $\widetilde{\bm{z}}$ with covariance $\widetilde{\Sigma}$, using a tilde to distinguish from the mean and covariance of the generative model. This generalizes the derivation in (buckley_free_2017, ) to the multivariate case, i.e.,
@@
For convenience, we label $a_{i}=z_{i}-\widetilde{z}_{i}$ and $A=\widetilde{\Sigma}^{-1}$, and $Z=\sqrt{(2\pi)^{n_{z}}\operatorname{det}A^{-1}}$. Substituting this ansatz into (A.17), we have
@@
where $D$ defined as before. This integral does not have a full closed-form solution due to the nonlinearity $f(W\bm{z})$. In the absence of a specific expression for this function, there are several ways to proceed. If $f(W\bm{z})$ admits a polynomial expansion in $\bm{z}$, then all odd terms in the integrand vanish, and the resulting integrals of the form $\int\!\mathrm{d}\bm{z}\,{\bm{z}}^{n}\,q(\bm{z})$ for $0<n\in 2\mathbb{Z}_{+}$ are easily done. A more abstract way of proceeding would be to perform an Edgeworth expansion of $p(\bm{x},\bm{z})$, with some assumption about how the corresponding cumulants integrate against a Gaussian. Intermediate between these two is the approach taken by (buckley_free_2017, ), which keeps $E$ unspecified, and formally Taylor expands it to second order around the mean of the variational posterior $\bm{z}=\widetilde{\bm{z}}$. However, this implicitly assumes that $f(W\bm{z})$ does not grow too fast in $\bm{z}$ (i.e., the polynomial expansion quickly truncates); otherwise, the fall-off of the Gaussian $q(\bm{z})$ is insufficiently fast for this approximation to apply for general functions. Furthermore, even in the special case where $\langle E(\bm{x},\bm{z})\rangle_{z}=\widetilde{\bm{z}}$, there is no guarantee that expanding the energy around this point will yield a good approximation to the integral. If $E(\bm{x},\bm{z})$ is sharply peaked at $\widetilde{\bm{z}}$ however, then we may proceed as in (buckley_free_2017, ); to second order,
@@
where we used the symmetry of the Hessian. This is the multivariate counterpart to eq. (17) in (buckley_free_2017, ). In sum, under a second order Taylor approximation of $E$, the optimal covariance of the variational is equal to the Hessian of $E$. Interestingly, including this term was shown to have practical consequences in (zahid_curvature-sensitive_2023, ) training generative PCNs, with substantially improved sample qualities and log-likelihoods as a result.
Appendix B Complementary Proofs
@@ B.2. Equivalence of FNNs and Discriminative PCNs During Testing
This appendix reproduces the proof in (whittington_approximation_2017, ; song_predictive_2021, ) that inference (testing) in FNNs and discriminative PCNs is equivalent, cf. section 2.2.5 in the main text. During testing, clamp $\bm{a}^{0}=\bm{x}^{(n)}$. We want to find
@@ Appendix C Computational Complexity
The results are shown in tables 8 for BP and 9 for IL and incremental IL. For both, the complexity without parallelization is shown, along with the results for parallelized layers ([PL]). It can be seen that with parallelized layers, standard IL is faster only if $T<L$, whereas incremental IL is faster by a factor $L$. For incremental IL, it should be noted that although the time complexity per weight update no longer scales with $T$, implementations so far suggest that, for a given epoch, one should still spend several inference iterations (i.e. E-steps) on a given datapoint/batch (in contrast with BP, which always performs a single forward and backward pass per batch). This means the total complexity would increase again with this factor. At the same time, it has been shown that low values $T$ (e.g., $T=3$ or $5$ (salvatori_stable_2023, ; alonso_understanding_2023)) appear to be sufficient, implying a substantial speed-up when $L$ is large (which is common in modern deep learning). However, it is still unknown what values of $T$ are appropriate in general.

### v2

@@ A.2.1. General Idea
$\equiv\int\prod_{\ell=1}^{L}d\bm{z}^{\ell}q(\{\bm{z}^{\ell}\})\ln\frac{q(\{\bm{z}^{\ell}\})}{p_{\theta}(\bm{x},\{\bm{z}^{\ell}\})}$
$=\underbrace{\int\prod_{\ell=1}^{L}\mathrm{d}\bm{z}^{\ell}q(\bar{\bm{z}})\ln q(\bar{\bm{z}})}_{H[\bar{q}]}-\underbrace{\int\prod_{\ell=1}^{L}\mathrm{d}\bm{z}^{\ell}q(\bar{\bm{z}})E(\bm{x},\bar{\bm{z}})}_{E_{\bar{q}}}~{},$
@@ A.2.2. PCNs of Rao & Ballard: Delta-function variational
Recent work (Millidge et al., 2022b) presents an attempted derivation of PC using a delta function as the variational posterior, in which it is claimed that the VFE simply becomes equal to the energy. Unfortunately however, this derivation contains a basic mathematical error, which we explain below. An alternative derivation (Buckley et al., 2017) also exists that instead uses a Gaussian as variational posterior, approximated using a Taylor series. Under certain assumptions, this leads to the claimed result. We review the both here and in the next subsection.
@@
In (Millidge et al., 2022a) this is erroneously taken to be zero. We speculate that this is based on the intuition that the “entropy” of a deterministic function like $\delta(z)$ should vanish. In fact however, the differential entropy of the Dirac delta function is $\infty$- ∞. If however one treats this as an additive constant to be dropped, then we have only the $E_{q}$ term, so that
@@
For this reason, (Millidge et al., 2022a) identify the VFE as simply the energy (or negative complete data log-likelihood) in this case.
@@ A.2.3. PCNs of Rao & Ballard: Gaussian variational
Now we instead take $q$ to be a multivariate Gaussian centered around $\widetilde{\bm{z}}$ with covariance $\widetilde{\Sigma}$, using a tilde to distinguish from the mean and covariance of the generative model. This generalizes the derivation in (Buckley et al., 2017) to the multivariate case, i.e.,
@@
For convenience, we label $a_{i}=z_{i}-\widetilde{z}_{i}$, $A=\widetilde{\Sigma}^{-1}$, and $Z=\sqrt{(2\pi)^{n_{z}}\operatorname{det}A^{-1}}$. Substituting this ansatz into (A.17), we have
@@
where $D=\frac{1}{2}\left((n_{x}+n_{z})\ln 2\pi+\ln\left[\det\Sigma_{x}\det\Sigma_{z}\right]\right)$. This integral does not have a full closed-form solution due to the nonlinearity $f(W\bm{z})$. In the absence of a specific expression for this function, there are several ways to proceed. If $f(W\bm{z})$ admits a polynomial expansion in $\bm{z}$, then all odd terms in the integrand vanish, and the resulting integrals of the form $\int\!\mathrm{d}\bm{z}\,{\bm{z}}^{n}\,q(\bm{z})$ for $0<n\in 2\mathbb{Z}_{+}$ are easily done. A more abstract way of proceeding would be to perform an Edgeworth expansion of $p(\bm{x},\bm{z})$, with some assumption about how the corresponding cumulants integrate against a Gaussian. Intermediate between these two is the approach taken by (Buckley et al., 2017), which keeps $E$ unspecified, and formally Taylor expands it to second order around the mean of the variational posterior $\bm{z}=\widetilde{\bm{z}}$. However, this implicitly assumes that $f(W\bm{z})$ does not grow too fast in $\bm{z}$ (i.e., the polynomial expansion quickly truncates); otherwise, the fall-off of the Gaussian $q(\bm{z})$ is insufficiently fast for this approximation to apply for general functions. Furthermore, even in the special case where $\langle E(\bm{x},\bm{z})\rangle_{z}=\widetilde{\bm{z}}$, there is no guarantee that expanding the energy around this point will yield a good approximation to the integral. If $E(\bm{x},\bm{z})$ is sharply peaked at $\widetilde{\bm{z}}$ however, then we may proceed as in (Buckley et al., 2017); to second order,
@@
where we used the symmetry of the Hessian. This is the multivariate counterpart to eq. (17) in (Buckley et al., 2017). In sum, under a second order Taylor approximation of $E$, the optimal covariance of the variational is equal to the Hessian of $E$. Interestingly, including this term was shown to have practical consequences in (Zahid et al., 2023b) training generative PCNs, with substantially improved sample qualities and log-likelihoods as a result. Hence, further studying the inclusion of higher-order terms in the minimization objective could potentially be an interesting avenue for further work. However, as noted by (Zahid et al., 2023b), such efforts need to balance potential increases in performance with increased computation. Moreover, better variational ansätze may encounter other practical issues, cf. (MacKay, 2003, 2001).
Appendix B Complementary Proofs and Miscellaneous
@@ B.2. Equivalence of FNNs and Discriminative PCNs During Testing
This appendix reproduces the proof in (Whittington and Bogacz, 2017; Song, 2021) that inference (testing) in FNNs and discriminative PCNs is equivalent, cf. section 2.2.4 in the main text. During testing, clamp $\bm{a}^{0}=\bm{x}^{(n)}$. We want to find
@@
B.3. Structure in PCNs
Here, we briefly expand on the discussion in section 2.3.3. What precisely defines the structure of a PCN? PCNs adds error nodes through (8) on top of activity nodes; but to get connections between layers like in FNNs, one should also include predictions, (7). However, illustrations like fig. 5 also include feedback connections, a key aspect of PCNs – hence they could additionally be included under structure. Mathematically, one way of defining these connections is by the activity rule (10). However, this equation is derived from the minimum of the objective function, which also defines the learning algorithm. This illustrates how structure and learning algorithm are more difficult to separate in PCNs.
This discussion explains why the term predictive coding in the literature sometimes appears to relate to a learning algorithm, and sometimes to different network architectures. Indeed, we have explained that it is in fact both. Nonetheless, some aspects of these algorithms are clearly ‘structural’ in nature: hence we keep using these terms when useful (as e.g. in fig. 2).
B.4. Z-IL pseudocode
Cf. section 4.1 in the main text. Note that to write down Z-IL most simply we here use notation that contradicts other sections. We use $\bm{\mu}^{\ell}=f(\bm{w}^{\ell+1}\bm{a}^{\ell+1})$, with data $\bm{x}^{(n)}$ clamped to $\bm{a}^{L}$ to and $\bm{y}^{(n)}$ to $\bm{a}^{0}$.
1:: $\bm{\mu}=f(\bm{w}^{\ell+1}\bm{a}^{\ell+1})$ // Downward predictions
2:: $\bm{a}^{L}=\bm{x}^{(n)}$, $\bm{a}^{0}=\bm{y}^{(n)}$ // Clamp data
4:: $\gamma=1$
5:for $t=0$ to $L$ do // Note: $T=L$
6:   for each $\ell$ do
9:   end for
10:end for
@@ Appendix C Computational Complexity
The results are shown in tables 7 for BP and 8 for IL and incremental IL. For both, the complexity without parallelization is shown, along with the results for parallelized layers ([PL]). It can be seen that with parallelized layers, standard IL is faster only if $T<L$, whereas incremental IL is faster by a factor $L$. For incremental IL, it should be noted that although the time complexity per weight update no longer scales with $T$, implementations so far suggest that, for a given epoch, one should still spend several inference iterations (i.e. E-steps) on a given datapoint/batch (in contrast with BP, which always performs a single forward and backward pass per batch). This means the total complexity would increase again with this factor. At the same time, it has been shown that low values $T$ (e.g., $T=3$ or $5$ (Salvatori et al., 2024b; Alonso et al., 2024)) appear to be sufficient, implying a substantial speed-up when $L$ is large (which is common in modern deep learning). However, it is still unknown what values of $T$ are appropriate in general.